# Bistability of travelling waves and wave-pinning states in a mass-conserved reaction-diffusion system:

### v3

@@ Abstract
Eukaryotic cells demonstrate a wide variety of dynamic patterns of filamentous actin (F-actin) and its regulators. Some of these patterns play important roles in cell functions, such as distinct motility modes, which motivate this study. We devise a mass-conserved reaction-diffusion model for active and inactive Rho-GTPase and F-actin in the cell cortex. The mass-conserved Rho-GTPase system promotes F-actin, which feeds back to inactivate the former. We study the model on a 1D periodic domain (edge of thin sheet-like cell) using bifurcation theory in the framework of spatial dynamics, complemented with numerical simulations. Among several discussed bifurcations, the analysis centers on the study of the codimension-2 long wavelength and finite wavenumber Hopf instability, in which we describe a rich structure of steady wave-pinning states (a.k.a. mesas, obeying the Maxwell construction), propagating coherent solutions (fronts and excitable pulses), and travelling and standing waves, all distinguished by mass conservation regimes and classified by domain sizes. Specifically, we highlight the unexpected conditions for bistability between steady wave-pinning and travelling wave states on moderate domain sizes. These results uncover and exemplify possible mechanisms of coexistence, robustness, and transitions between distinct cellular motility modes, including directed migration, turning, and ruffling. More broadly, the results indicate that dissipative reaction-diffusion models comprising mass conservation have distinct pattern formation mechanisms that motivate further investigations, such as the unfolding of codimension-3 instabilities and T-points.
@@ 1 Introduction
Based on previous works [17, 20, 39, 42, 98] that show distinct regimes of patterns in parameter space, is it possible to identify a mechanism for bistability of travelling waves and polarization in dissipative reaction-diffusion models with mass conservation?
@@
To address the above, we simplify a three-variable mass-conserving reaction-diffusion model by Holmes et al. [38], in one space dimension (1D) on a ring (periodic cell edge) and use bifurcation analysis to dissect it.
@@
modification of the dissipative mass-conserving reaction-diffusion model for actin waves by Holmes et al. [38];
@@ 2 Model equations for the GTPase-actin system
The starting point of our modelling is the two-component mass-conserved reaction-diffusion system for cell polarization proposed by Mori et al. [63] (Figure 2(a)). Essential features of this system are slow-fast rates of diffusion for active ($u$) and inactive ($v$) GTPase, respectively, and nonlinear kinetic terms that admit bistability in some parameter regimes [34]. On the timescales of interest (seconds, minutes) the GTPase is neither produced nor degraded, so Mori et al. [63] assumed the total mass of GTPase was conserved. Therefore, the bistability differs from that of generic dissipative models, as there is a single nullcline and the flow is restricted by mass conservation. In particular, since the system can be written in a gradient form, the steady-states depend on the Maxwell construction and are non-hyperbolic [10, 16, 21, 49, 64, 69, 75]. This system produces a spatially nonuniform steady solution (denoted a wave-pinning solution) associated with cell polarity: a high plateau of active GTPase (corresponding to the leading edge of a cell) separated by an interface from a low active GTPase plateau. Note that wave-pinning (WP) solutions are also known as “mesa” states in the context of porous media [28] and chemical patterning [48]. Holmes et al. [38] extended the Mori et al. model [63] to capture “actin waves” in reasoning that is similar to the idea behind the classic FitzHugh-Nagumo (FHN) system [30, 67], where a bistable system is coupled to slow negative feedback: the $(u,v)$-system acts as the “bistable” part, and slow negative feedback is provided by F-actin ($F$), whose role is similar to a “refractory variable” in the FHN model (Figure 2(b)). The experiments of Michaud et al. [57] describe a real cell system with interactions as in the Holmes et al. model (although the authors chose to rederive their own, rather similar, model system): the GTPase is Rho, it self-activates (via the GEF Ect2), and F-actin recruits a GAP (RGA-3/4) that inactivates Rho. In what follows, we simplify the original WP and actin waves models to produce the model that we will analyze in the remainder of this study.
@@ 3 Linear analysis of homogeneous steady states (HSSs)
In our setting, we are mainly interested in two instabilities: the steady long wavelength (LW) and the finite wavenumber Hopf (WB). The LW occurs as in gradient systems, i.e., this instability occurs when the concavity of $\max_{\sigma}\text{Re}(\sigma(q))$ changes sign at $q=0$ [23], since at least one $\sigma(0)=0$ by mass conservation [9]. The WB is an oscillatory type, simultaneously giving rise to both travelling waves (TW) and standing waves (SWs) [45]. In addition, we find it informative to monitor bifurcations occurring when the HSS is linearly unstable, i.e., not only the LW and the WB but also the homogenous Hopf (HB). Consequently, we show in Figure 3, selected two-parameter diagrams while varying $(s,b)$, for $M=2$ and $M=4.5$, in which we identify saddle-nodes (SNs) of HSS (black) and the locus of bifurcation onsets: HB (red), WB (blue), and LW (green). For an illustration of the bifurcation onsets and the following analysis, we focus on a selected slice at $b=b_{c}\approx 0.067$ (dotted line in Figure 3(a)) exhibiting a codimension-2 point of LW/WB instabilities for $M=2$, as shown in Figure 4. Here we have two LW bifurcations, where the second is near the $SN_{Q_{0/1}^{\ast}}$. In Appendix A, we show additional slices where more codimension-2 points exist (see Figs. 20, 21, 22, and 23). We focus on $M=2$ since for $M=4.5$, the LW loop is the most inner one (see Figure 3(b)), excluding the occurrence of the LW/WB instability. Nevertheless, we briefly demonstrate the similarities and differences to the case of $M=4.5$ (in Section 5) and compare with the results obtained by Yochelis et al. [98], see also Appendix B. We note that for $b\approx 0.14$, in a region above the cusp bifurcation of HSSs (shaded yellow), there is another codimension-2 LW/WB bifurcation. The latter is outside the scope of this work and will be addressed elsewhere.
@@ 4 Bifurcation criticality of single-wavelength oscillatory solutions
We proceed here with the analysis of the bifurcation criticality and relative linear stability of oscillatory solutions emerging from the LW/WB instability, $s=s_{WB}=s_{c}^{t}\approx 0.409$ and $b=b_{c}\approx 0.067$, up to larger values of $b$ and about $s=0.6$, see Figure 3(a). At the LW/WB onset, the steady solutions bifurcate subcritically (see Figure 9), namely toward the linearly stable regime, while the travelling ($\text{TW}_{\lambda}$) and standing ($\text{SW}_{\lambda}$) waves with critical wavelength $\lambda=2\pi/q_{c}\approx 3.093$ are supercritical with $\omega_{c}\approx 0.388$ [40]. To check if the criticalities of the single-wavelength TWs and SWs remain supercritical, we perform weakly nonlinear stability analysis, a.k.a. employing multiple timescales or the amplitude equation method. We ignore weak spatial modulations, including the impact of the large-scale mode [56, 91], due to the relationship to cells with perimeters comparable to a small number of spatial wavelengths.
Let $\epsilon$ be a small parameter such that $\epsilon^{2}\propto|s-s_{WB}|\ll 1$ is a measure of the system’s distance from the onset of instability. Then $s=s_{WB}+\epsilon^{2}\widetilde{s}$, where $\widetilde{s}$ is of $\mathcal{O}\left(1\right)$ and respectively, $Q^{\ast}=Q_{c}^{\ast}(s_{WB})+\epsilon^{2}\widetilde{Q}^{\ast}$. We set the power of $\epsilon$ in the slow temporal scale by observing the dependence of the growth rate $\sigma$ on $\epsilon$. Finding a quadratic dependence, $\sigma\propto\epsilon^{2}$, we arrive at the temporal scaling $\tau=\epsilon^{2}t$. The scaling for the slowly varying amplitudes for the left/right propagating waves is, therefore, set to $\widetilde{\textbf{\em B}}_{L}(\tau)=\epsilon\textbf{\em B}_{L}$, $\widetilde{\textbf{\em B}}_{R}(\tau)=\epsilon\textbf{\em B}_{R}$, where the respective derived equations read (for details, see Appendix C):
@@
Figure 5: (a) Bifurcating branches of the primary travelling and standing waves (with wavelength $\lambda\approx 3.09$) obtained at the codimension-2 onset at $s=s_{WB}=s_{c}^{t}\approx 0.409$ and $b=b_{c}\approx 0.067$ (see the slice in Figure 3(a)), hereafter $\text{TW}_{\lambda}$ and $\text{SW}_{\lambda}$, respectively. The solutions are projected onto the maximum F-actin concentration, $F_{\max}$. We compare results from the amplitude equations Eq. 18 and Eq. 19 to numerical continuation, over spatial wavelength $\lambda$ and, in the case of SW, also over the temporal period $T$ (also obtained at the onset): Linearly stable TWs (from continuation: solid blue; from amplitude equations: dashed-dotted blue) and linearly unstable SWs (from continuation: dashed orange; from amplitude equations: dotted orange). The inset shows the difference between the total amplitudes of the TWs and SWs according to Eq. 19 ($\Delta\tilde{\rho}={\widetilde{\rho}}_{TW}-{\widetilde{\rho}}_{SW}$), indicating that TWs are larger than SWs. (b) Same as (a) but with TWs being projected onto the phase wave speed $c$ (i.e., the initial propagation speed, left vertical axis) and SWs onto the time-period $T$ (right vertical axis), see Eq. 22 for the corresponding expressions from the amplitude equations. Coefficients for the amplitude equations are obtained numerically using Mathematica: $\alpha=2.13528-1.66874i$, $\gamma=0.06019+0.01434i$, $\eta=0.06500+0.16646i$, where $\widetilde{a}_{re}=\gamma_{re}-\eta_{re}\approx-0.00482$, $\widetilde{b}_{re}=-\gamma_{re}\approx-0.06019$, $\widetilde{a}_{im}=\gamma_{im}-\eta_{im}\approx-0.15211$, and $\widetilde{b}_{im}=-\gamma_{im}\approx-0.01434$. Panels (c,d) provide the same comparison but with $b=0.25$ and the finite wavenumber Hopf onset occurring at $s=s_{WB}\approx 0.522$ with a critical wavelength $\lambda\approx 3.13$, and respective coefficients: $\alpha=1.26708-0.46261i$, $\gamma=0.14570-0.12758i$, $\eta=0.38413-0.11440i$, where $\widetilde{a}_{re}=\gamma_{re}-\eta_{re}\approx-0.23843$, $\widetilde{b}_{re}=-\gamma_{re}\approx-0.14570$, $\widetilde{a}_{im}=\gamma_{im}-\eta_{im}\approx-0.01318$, and $\widetilde{b}_{im}=-\gamma_{im}\approx 0.12758$. Other parameters as in Table 1 with $M=2$.
@@
Analysis of these equations is convenient using polar forms $B=\rho\exp{i\Phi}$ with the subscripts $re/im$ denoting the real/imaginary parts of coefficients, and for convenience, rewrite in a canonic form [45]
(16a) $\dot{\rho}_{L}$ $=\rho_{L}\left[\widetilde{\alpha}_{re}\left(s-s_{WB}\right)+\widetilde{a}_{re}\rho_{R}^{2}+\widetilde{b}_{re}\left(\rho_{L}^{2}+\rho_{R}^{2}\right)\right],$
(16b) $\dot{\rho}_{R}$ $=\rho_{R}\left[\widetilde{\alpha}_{re}\left(s-s_{WB}\right)+\widetilde{a}_{re}\rho_{L}^{2}+\widetilde{b}_{re}\left(\rho_{L}^{2}+\rho_{R}^{2}\right)\right],$
(16c) $\dot{\Phi}_{L}$ $=\widetilde{\alpha}_{im}\left(s-s_{WB}\right)+\widetilde{a}_{im}\rho_{R}^{2}+\widetilde{b}_{im}\left(\rho_{L}^{2}+\rho_{R}^{2}\right),$
(16d) $\dot{\Phi}_{R}$ $=-\widetilde{\alpha}_{im}\left(s-s_{WB}\right)-\widetilde{a}_{im}\rho_{L}^{2}-\widetilde{b}_{im}\left(\rho_{L}^{2}+\rho_{R}^{2}\right).$
Steady solutions to Eq. 16 are
@@
(17a) $\rho$ $=\sqrt{\frac{\widetilde{\alpha}_{re}}{\widetilde{b}_{re}}\left(s_{WB}-s\right)},\quad{\rm for\,\,TWs\,\,with\,\,}(\rho_{L},0)=(\rho,0){\rm\,\,or\,\,}(0,\rho_{R})=(0,\rho),$
@@
(17b) $\rho$ $=\sqrt{\frac{\widetilde{\alpha}_{re}}{\widetilde{a}_{re}+2\widetilde{b}_{re}}\left(s_{WB}-s\right)},\quad{\rm for\,\,SWs\,\,with\,\,}(\rho_{L},\rho_{R})=(\rho,\rho),$
@@
(19) $\widetilde{\rho}=\sqrt{\rho_{L}^{2}+\rho_{R}^{2}}=\begin{cases}\sqrt{\dfrac{\widetilde{\alpha}_{re}}{\widetilde{b}_{re}}\left(s_{WB}-s\right)},\quad\rm for\,\,TWs;\\ \\ \sqrt{\dfrac{2\widetilde{\alpha}_{re}}{\widetilde{a}_{re}+2\widetilde{b}_{re}}\left(s_{WB}-s\right)},\quad\rm for\,\,SWs,\end{cases}$
@@
(20) $\Omega_{TW}$ $=\left(\widetilde{\alpha}_{im}-\widetilde{\alpha}_{re}\frac{\widetilde{b}_{im}}{\widetilde{b}_{re}}\right)\left(s-s_{WB}\right),$
(21) $\Omega_{SW}$ $=\left(\widetilde{\alpha}_{im}-\widetilde{\alpha}_{re}\frac{\widetilde{a}_{im}+2\widetilde{b}_{im}}{\widetilde{a}_{re}+2\widetilde{b}_{re}}\right)\left(s-s_{WB}\right),$
@@
Conducting the above computations for the left side of the WB loop, i.e., $b>b_{c}$, we find that the criticalities of TW and SW persist. As can be seen in Figure 5, the results derived from amplitude equations agree well far from the codimension-2 LW/WB instability (bottom panels), whereas near the codimension-2 LW/WB instability only qualitatively properties agree (top panels). This is because for $s>s^{t}_{c}$, the instability is for both the LW and the oscillatory modes so that the coupling between them affects the supercritical bifurcating oscillatory states. Interestingly, the speed of TWs near the codimension-2 LW/WB instability first decreases and then increases as opposed to only increasing in the absence of the LW onset, as shown in Figure 5. We also note that below the codimension-2 point, i.e., $b<b_{c}$, both TW and SW are initially unstable since there, WB lies inside the LW loop (i.e. linearly unstable regime of $Q_{2}^{\ast}$). Nevertheless, TWs of critical period or otherwise may still gain stability through other mechanisms far from the onset. Some of these will be demonstrated in the following section.

### v4

@@ Abstract
Eukaryotic cells demonstrate a wide variety of dynamic patterns of filamentous actin (F-actin) and its regulators. Some of these patterns play important roles in cell functions, such as distinct motility modes, which motivate this study. We devise a mass-conserved reaction-diffusion model for active and inactive Rho-GTPase and F-actin in the cell cortex. The mass-conserved Rho-GTPase system promotes F-actin, which feeds back to inactivate the former. We study the model on a 1D periodic domain (edge of thin sheet-like cell) using bifurcation theory in the framework of spatial dynamics, complemented with numerical simulations. Among several discussed bifurcations, the analysis centers on the study of the codimension-2 long wavelength and finite wavenumber Hopf instability, in which we describe a rich structure of steady wave-pinning states (a.k.a. mesas, obeying the Maxwell construction), propagating coherent solutions (fronts and excitable pulses), and travelling and standing waves, all distinguished by mass conservation regimes and classified by domain sizes. Specifically, we highlight the unexpected conditions for bistability between steady wave-pinning and travelling wave states on moderate domain sizes, i.e., unfolding through domain length. These results uncover and exemplify possible mechanisms of coexistence, robustness, and transitions between distinct cellular motility modes, including directed migration, turning, and ruffling. More broadly, the results indicate that non-gradient reaction-diffusion models comprising mass conservation have distinct pattern formation mechanisms that motivate further investigations, such as the unfolding of codimension-3 instabilities and T-points.
@@ 1 Introduction
Based on previous works [17, 20, 39, 42, 99] that show distinct regimes of patterns in parameter space, is it possible to identify a mechanism for bistability of travelling waves and polarization in non-gradient reaction-diffusion models with mass conservation?
@@
To address the above, we modify a three-variable mass-conserving reaction-diffusion model by Holmes et al. [38], in one space dimension (1D) on a ring (periodic cell edge) and use bifurcation analysis to dissect it.
@@
modification of the non-gradient mass-conserving reaction-diffusion model for actin waves by Holmes et al. [38], by replacing Hill functions with polynomial nonlinearities representing the coupling between the GTPase and F-actin;
@@ 2 Model equations for the GTPase-actin system
The starting point of our modelling is the two-component mass-conserved reaction-diffusion system for cell polarization proposed by Mori et al. [63] (Figure 2(a)). Essential features of this system are slow-fast rates of diffusion for active ($u$) and inactive ($v$) GTPase, respectively, and nonlinear kinetic terms that admit bistability in some parameter regimes [34]. On the timescales of interest (seconds, minutes) the GTPase is neither produced nor degraded, so Mori et al. [63] assumed the total mass of GTPase was conserved. Therefore, the bistability differs from that of generic non-gradient models, as there is a single nullcline and the flow is restricted by mass conservation. In particular, since the system can be written in a gradient form, the steady-states depend on the Maxwell construction and are non-hyperbolic [10, 16, 21, 49, 64, 69, 76]. This system produces a spatially nonuniform steady solution (denoted a wave-pinning solution) associated with cell polarity: a high plateau of active GTPase (corresponding to the leading edge of a cell) separated by an interface from a low active GTPase plateau. Note that wave-pinning (WP) solutions are also known as “mesa” states in the context of porous media [28] and chemical patterning [48]. Holmes et al. [38] extended the Mori et al. model [63] to capture “actin waves” in reasoning that is similar to the idea behind the classic FitzHugh-Nagumo (FHN) system [30, 67], where a bistable system is coupled to slow negative feedback: the $(u,v)$-system acts as the “bistable” part, and slow negative feedback is provided by F-actin ($F$), whose role is similar to a “refractory variable” in the FHN model (Figure 2(b)). The experiments of Michaud et al. [57] describe a real cell system with interactions as in the Holmes et al. model (although the authors chose to rederive their own, rather similar, model system): the GTPase is Rho, it self-activates (via the GEF Ect2), and F-actin recruits a GAP (RGA-3/4) that inactivates Rho. In what follows, we simplify the original WP and actin waves models to produce the model that we will analyze in the remainder of this study.
@@ 3 Linear analysis of homogeneous steady states (HSSs)
In our setting, we are mainly interested in two instabilities: the steady long wavelength (LW) and the finite wavenumber Hopf (WB). The LW occurs as in gradient systems, i.e., this instability occurs when the curvature of $\max_{\sigma}\text{Re}(\sigma(q))$ changes sign at $q=0$ [23], since at least one $\sigma(0)=0$ by mass conservation [9]. The WB is an oscillatory type, simultaneously giving rise to both travelling waves (TW) and standing waves (SWs) [45]. In addition, we find it informative to monitor bifurcations occurring when the HSS is linearly unstable, i.e., not only the LW and the WB but also the homogenous Hopf (HB). Consequently, we show in Figure 3, selected two-parameter diagrams while varying $(s,b)$, for $M=2$ and $M=4.5$, in which we identify saddle-nodes (SNs) of HSS (black) and the locus of bifurcation onsets: HB (red), WB (blue), and LW (green). For an illustration of the bifurcation onsets and the following analysis, we focus on a selected slice at $b=b_{c}\approx 0.067$ (dotted line in Figure 3(a)) exhibiting a codimension-2 point of LW/WB instabilities for $M=2$, as shown in Figure 4. Here we have two LW bifurcations, where the second is near the $SN_{Q_{0/1}^{\ast}}$. In Appendix A, we show additional slices where more codimension-2 points exist (see Figs. 20, 21, 22, and 23). We focus on $M=2$ since for $M=4.5$, the LW loop is the most inner one (see Figure 3(b)), excluding the occurrence of the LW/WB instability. Nevertheless, we briefly demonstrate the similarities and differences to the case of $M=4.5$ (in Section 5) and compare with the results obtained by Yochelis et al. [99], see also Appendix B. We note that for $b\approx 0.14$, in a region above the cusp bifurcation of HSSs (shaded yellow), there is another codimension-2 LW/WB bifurcation. The latter is outside the scope of this work and will be addressed elsewhere.
@@ 4 Bifurcation criticality of single-wavelength oscillatory solutions
We proceed here with the analysis of the bifurcation criticality and relative linear stability of oscillatory solutions emerging from the LW/WB instability, $s=s_{WB}=s_{c}^{t}\approx 0.409$ and $b=b_{c}\approx 0.067$, up to larger values of $b$ and about $s=0.6$, see Figure 3(a). At the LW/WB onset, the steady solutions bifurcate subcritically (see Figure 9), namely toward the linearly stable regime, while the travelling ($\text{TW}_{\lambda}$) and standing ($\text{SW}_{\lambda}$) waves with critical wavelength $\lambda=2\pi/q_{c}\approx 3.093$ are supercritical with $\omega_{c}\approx 0.388$ [40]. To check if the criticalities of the single-wavelength TWs and SWs remain supercritical for the varied parameter values, we perform weakly nonlinear stability analysis, a.k.a. employing multiple timescales or the amplitude equation method. We ignore weak spatial modulations, including the impact of the large-scale mode [56, 92], due to the relationship to cells with perimeters comparable to a small number of spatial wavelengths.
Let $\epsilon$ be a small parameter such that $\epsilon^{2}\propto|s-s_{WB}|\ll 1$ is a measure of the system’s distance from the onset of instability. Then $s=s_{WB}+\epsilon^{2}\widetilde{s}$, where $\widetilde{s}$ is of $\mathcal{O}\left(1\right)$ and respectively, $Q^{\ast}=Q_{c}^{\ast}(s_{WB})+\epsilon^{2}\widetilde{Q}^{\ast}$. We set the power of $\epsilon$ in the slow temporal scale by observing the dependence of the growth rate $\sigma$ on $\epsilon$. Finding a quadratic dependence, $\sigma\propto\epsilon^{2}$, we arrive at the temporal scaling $\tau=\epsilon^{2}t$. The amplitudes for the left/right propagating waves (slowly varied in time), are obtained by expanding $Q$ in powers of $\epsilon$,
and using the ansatz:
(16c) $F_{1}$ $=B_{LF}(\tau)e^{i(\omega_{c}t+q_{c}x)}+B_{RF}(\tau)e^{i(\omega_{c}t-q_{c}x)}+c.c..$
In (16), $\omega_{c}$ and $q_{c}$ denote the critical frequency and wavenumber at the instability onset, and subscripts $L,R$ stand for left and right, respectively. Following a standard procedure (detailed in Appendix C), leads to the equations:
@@
Figure 5: (a) Bifurcating branches of the primary travelling and standing waves (with wavelength $\lambda\approx 3.09$) obtained at the codimension-2 onset at $s=s_{WB}=s_{c}^{t}\approx 0.409$ and $b=b_{c}\approx 0.067$ (see the slice in Figure 3(a)), hereafter $\text{TW}_{\lambda}$ and $\text{SW}_{\lambda}$, respectively. The solutions are projected onto the maximum F-actin concentration, $F_{\max}$. We compare results from the amplitude equations Eq. 20 and Eq. 21 to numerical continuation, over spatial wavelength $\lambda$ and, in the case of SW, also over the temporal period $T$ (also obtained at the onset): Linearly stable TWs (from continuation: solid blue; from amplitude equations: dashed-dotted blue) and linearly unstable SWs (from continuation: dashed orange; from amplitude equations: dotted orange). The inset shows the difference between the total amplitudes of the TWs and SWs according to Eq. 21 ($\Delta\tilde{\rho}={\widetilde{\rho}}_{TW}-{\widetilde{\rho}}_{SW}$), indicating that TWs are larger than SWs. (b) Same as (a) but with TWs being projected onto the phase wave speed $c$ (i.e., the initial propagation speed, left vertical axis) and SWs onto the time-period $T$ (right vertical axis), see Eq. 24 for the corresponding expressions from the amplitude equations. Coefficients for the amplitude equations are obtained numerically using Mathematica: $\alpha=2.13528-1.66874i$, $\gamma=0.06019+0.01434i$, $\eta=0.06500+0.16646i$, $\widehat{a}\approx-0.00482-0.15211i$, and $\widehat{b}\approx-0.06019-0.01434i$. Panels (c,d) provide the same comparison but at $b=0.25$, where the finite wavenumber Hopf onset occurring at $s=s_{WB}\approx 0.522$ with a critical wavelength $\lambda\approx 3.13$, and respective coefficients: $\alpha=1.26708-0.46261i$, $\gamma=0.14570-0.12758i$, $\eta=0.38413-0.11440i$, $\widehat{a}\approx-0.23843-0.01318i$, and $\widehat{b}\approx-0.14570+0.12758i$. Other parameters as in Table 1 with $M=2$.
@@
Analysis of these equations is convenient using polar forms $B=\rho\exp{i\Phi}$ with the subscripts $re/im$ denoting the real/imaginary parts of coefficients, and for convenience, rewrite in a canonic form of total amplitudes [45]
(18a) $\dot{\rho}_{L}$ $=\rho_{L}\left[\alpha_{re}\left(s-s_{WB}\right)+\widehat{a}_{re}\rho_{R}^{2}+\widehat{b}_{re}\left(\rho_{L}^{2}+\rho_{R}^{2}\right)\right],$
(18b) $\dot{\rho}_{R}$ $=\rho_{R}\left[\alpha_{re}\left(s-s_{WB}\right)+\widehat{a}_{re}\rho_{L}^{2}+\widehat{b}_{re}\left(\rho_{L}^{2}+\rho_{R}^{2}\right)\right],$
(18c) $\dot{\Phi}_{L}$ $=\alpha_{im}\left(s-s_{WB}\right)+\widehat{a}_{im}\rho_{R}^{2}+\widehat{b}_{im}\left(\rho_{L}^{2}+\rho_{R}^{2}\right),$
(18d) $\dot{\Phi}_{R}$ $=-\alpha_{im}\left(s-s_{WB}\right)-\widehat{a}_{im}\rho_{L}^{2}-\widehat{b}_{im}\left(\rho_{L}^{2}+\rho_{R}^{2}\right),$
where $\widehat{a}_{re}\equiv\gamma_{re}-\eta_{re}$, $\widehat{a}_{im}\equiv\gamma_{im}-\eta_{im}$, and $\widehat{b}\equiv-\gamma$. Steady solutions to Eq. 18 are
@@
(19a) $\rho$ $=\sqrt{\frac{\alpha_{re}}{\widehat{b}_{re}}\left(s_{WB}-s\right)},\quad{\rm for\,\,TWs\,\,with\,\,}(\rho_{L},0)=(\rho,0){\rm\,\,or\,\,}(0,\rho_{R})=(0,\rho),$
@@
(19b) $\rho$ $=\sqrt{\frac{\alpha_{re}}{\widehat{a}_{re}+2\widehat{b}_{re}}\left(s_{WB}-s\right)},\quad{\rm for\,\,SWs\,\,with\,\,}(\rho_{L},\rho_{R})=(\rho,\rho),$
@@
(21) $\widetilde{\rho}=\sqrt{\rho_{L}^{2}+\rho_{R}^{2}}=\begin{cases}\sqrt{\dfrac{\alpha_{re}}{\widehat{b}_{re}}\left(s_{WB}-s\right)},\quad\rm for\,\,TWs;\\ \\ \sqrt{\dfrac{2\alpha_{re}}{\widehat{a}_{re}+2\widehat{b}_{re}}\left(s_{WB}-s\right)},\quad\rm for\,\,SWs,\end{cases}$
@@
(22) $\Omega_{TW}$ $=\left(\alpha_{im}-\alpha_{re}\frac{\widehat{b}_{im}}{\widehat{b}_{re}}\right)\left(s-s_{WB}\right),$
(23) $\Omega_{SW}$ $=\left(\alpha_{im}-\alpha_{re}\frac{\widehat{a}_{im}+2\widehat{b}_{im}}{\widehat{a}_{re}+2\widehat{b}_{re}}\right)\left(s-s_{WB}\right),$
@@
Conducting the above computations for the left side of the WB loop, i.e., $b>b_{c}$, we find that the criticalities of TW and SW persist. As can be seen in Figure 5, the results derived from amplitude equations agree well far from the codimension-2 LW/WB instability (bottom panels), whereas near the codimension-2 LW/WB instability only qualitatively properties agree (top panels). This is because for $s>s^{t}_{c}$, the instability is for both the LW and the oscillatory modes so that the coupling between them affects the supercritical bifurcating oscillatory states. Interestingly, the speed of TWs near the codimension-2 LW/WB instability first decreases and then increases as opposed to only increasing in the absence of the LW onset, as shown in Figure 5. Naturally, in the absence of spatial modulations, the derived amplitude equations cannot address this inconsistency. We also note that below the codimension-2 point, i.e., $b<b_{c}$, both TW and SW are initially unstable since there, WB lies inside the LW loop (i.e. linearly unstable regime of $Q_{2}^{\ast}$). Nevertheless, TWs of critical period or otherwise may still gain stability through other mechanisms far from the onset. Some of these will be demonstrated in the following section.